Solve y''+9y=f(t) with y(0)=0 and y'(0)=0 and f(t)=sin(t) when t \in [0, \pi] and zero otherwise

Use laplace transform
Answer #1

o tot Y + Gy = f(t) 17)-S Sint 05t< 910-0, 460)=0 f(t) = Sint- Lly + gy) = L( f(t)) 219 +9 hly) = 4( Sint - UCE-T) Sint)

L(Y(s)) = 41965.) - _ yle) sin lt) - 1 Sin(38) tft sin(-7) sen (314-))) UCE-1) 24 I dit) - Sinit) Ult- Sinlt) - Sin (3+) 8 २